# Estimating Heterogeneous Causal Effects of High-Dimensional Treatments:

### v4

@@ F.2.2 Updating Moderator Parameters
where $\pi_{k}(\bm{X}_{i})$ and $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ are functions of $\bm{\phi}_{k}$. Note that if $\gamma=0$, this simplifies to a multinomial logistic regression with $\{\mathbb{E}[\bm{1}\{Z_{i}=k\}]\}_{k=1}^{K}$ as the outcome. We perform the $M$-Step by using a few steps of a standard optimizer (e.g., L-BFGS) to increase $Q_{\phi}$ and thus obtain $\bm{\phi}^{(t+1)}$.
F.3 Degrees of Freedom
@@
F.4 Computational Improvements
While the algorithm above provides a valid way to locate a posterior mode, our estimation problem is complex and high-dimensional. Furthermore, given the complex posterior implied by mixture of experts models, we derived a number of computational strategies to improve convergence. We use the SQUAREM algorithm (Varadhan and Roland 2008) and a generalized EM algorithm to update $\bm{\beta}$ using a conjugate gradient approach and $\bm{\phi}$ using a few steps of L-BFGS.
@@ F.4 Computational Improvements
Supplementary Material G Extensions to the Basic Model
@@ G.1 Repeated Observations
G.2 Forced Choice Conjoint Design
@@ G.3 Standardization Weights
This has no implication on the rank of the stacked $\bm{F}_{g}$ (and thus the results in Appendix E) as they are all positive and thus only slightly modify the $E$-Step.
We employ weights in all of our analyses to account for the fact that different factors $j$ may have different number of levels $L_{j}$. We use a generalization of the weights in Bondell and Reich (2009) to the case of penalized differences. Specifically, consider the over-parameterized model in Appendix G.4 where the penalty can be written entirely on the differences $\bm{\delta}_{\mathrm{Main}}$, $\bm{\delta}_{\mathrm{Int}}$, $\bm{\delta}_{\mathrm{Main-Copy}}$. Note that each of those penalties has a simple (group) LASSO form and thus we adopt the approach in Lim and Hastie (2015) of weighting by the Frobenius norm of the associated columns in $\bm{T}_{\mathrm{LOG}}$, i.e. the over-parameterized design matrix. At slight abuse of notation, define $[\bm{T}_{\mathrm{LOG}}]_{g}$ as the columns of $\bm{T}_{\mathrm{LOG}}$ corresponding to the differences penalized in the (group) lasso $g$, the weight can be expressed as follows:
@@
G.4 Latent Overlapping Groups
@@
Supplementary Material H Estimators
@@
H.1 Factorial designs
H.1.1 Without restrictions on randomization
@@
H.1.2 With restrictions on randomization
@@
H.2 Conjoint designs
H.2.1 Without restrictions on randomization
@@
H.2.2 With restrictions on randomization
Similar to Appendix H.1.2, adjustments to estimation need to be made when we have restricted randomizations. We again will do this by dropping profiles that have levels of factors not allowable for all levels of the factor(s) whose effects we are estimating (e.g., profiles with “4th grade” for education when estimating an effect for profession). However, now we estimate the effect for the right profile and the effect for the left profile, and then average the two (they should be equal under symmetry). When estimating the effect for the right profile, therefore, we will only drop pairings if the right profile has a level that is not allowed for some level of the factor we are estimating an effect of. For example, dropping pairings where the right profile has “4th grade” as level of education when estimating main effects of profession because “doctor” cannot have level “4th grade.” Again, this will drop more profiles than those that are not allowed under randomization to ensure an “apples-to-apples” comparison across levels of profession.
@@ H.2.2 With restrictions on randomization
Supplementary Material I Quantification of Uncertainty
@@
$\mathcal{I}_{L}(\hat{\bm{\theta}})=E_{p\left(\{Z_{i}\}_{i=1}^{N}|\hat{\bm{\theta}}\right)}\left[-\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}\bm{\theta}^{\top}}\right]-\mathrm{Var}_{p\left(\{Z_{i}\}_{i=1}^{N}|\hat{\bm{\theta}}\right)}\left[\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}}\right]$ (A18)
@@ I.1 Derivation of Hessian
$\tilde{S}^{c}(\bm{\phi}_{k})$ $=\sum_{i=1}^{N}\left[Z_{ik}-\pi_{ik}\right]\bm{X}_{i}+\frac{\partial\log p(\{\bm{\phi}_{k}\})}{\partial\bm{\phi}_{k}}+\sum_{k^{\prime}=1}^{K}m\gamma\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}}-\lambda\gamma\bar{\pi}_{k^{\prime}}^{\gamma-1}\cdot\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\cdot\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]$
@@
$+\sum_{k^{\prime}=1}^{K}-\lambda\gamma\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]\left[I(\gamma\notin\{0,1\})\cdot(\gamma-1)\bar{\pi}_{k^{\prime}}^{\gamma-2}\cdot{\color[rgb]{0,0,0}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\right]}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{\ell}}\right]^{\top}+\bar{\pi}_{k^{\prime}}^{\gamma-1}\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}\right]$
@@ I.2 Repeated Observations
$\mathrm{Cov}\left[\tilde{S}^{c}(\mu),\tilde{S}^{c}(\mu)\right]$ $=\sum_{i=1}^{N}\left[\sum_{k=1}^{K}\sum_{k^{\prime}=1}^{K}\mathbb{E}(Z_{ik})\left(I[k=k^{\prime}]-\mathbb{E}(Z_{ik^{\prime}})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}^{\top}\right]\right]$
@@
I.3 Standard Errors on Other Quantities of Interest
Given the above results, we derive an approximate covariance matrix on $\hat{\bm{\theta}}$. We calculate uncertainty on other quantities of interest, e.g. AMCE and marginal effects, using the multivariate delta method. As almost all of our quantities of interest can be expressed as (weighted) sums or averages over individuals $i\in\{1,\cdots,N\}$, calculating the requisite gradient for the multivariate delta method simply requires calculating the relevant derivative for each observation. For example, all derivatives needed in the AMCE are of the following form; see Appendix H for more details.
@@ I.3 Standard Errors on Other Quantities of Interest
Supplementary Material J Simulations
@@
J.1 Setup
We generate the $\bm{\beta}_{k}$ used in our simulations following Equation 4 and calibrating their implied AMCEs to be roughly comparable to the magnitude found in our empirical example, i.e. ranging between around $-0.30$ and $0.30$. The $\bm{\beta}_{k}$ and $\{\bm{\phi}_{k}\}_{k=2}^{3}$ used in all simulations are determined using one draw from the following procedure:
@@
In terms of simulating the treatment profiles and outcome, for each individual $i$, we draw a group membership $Z_{i}$ using $\bm{\pi}_{i}$ generating using $\bm{X}_{i}$, $\bm{\phi}$ and Equation 4. For each task $t$, we then randomly draw a pair of treatments and then, given $Z_{i}$, draw the outcome $Y_{i}$ given their observed treatments using the model in the main text.
@@
To address this issue, we consider an alternative procedure based on sample splitting. We fit the model using half of the data (selected at random) and then refit the model. To refit the model, we hold fixed the sparsity pattern estimated in the original estimation hold (i.e., which levels are fused together) using a tolerance of $10^{-3}$. We also fix the estimated moderator relationship, i.e. $\pi_{k}(\bm{X}_{i})$, and only estimate the treatment effect coefficients after fusion. Algorithm A2 states the procedure. To calculate the average marginal effects, as noted in Appendix H, we use the empirical distribution of treatments to marginalize over other factors. In this split version, we also use the distribution from the full dataset.
@@
J.3.1 Different Choice of $K$
We first consider how different choices of $K$ impact our results in the simulation study. To do this, we focus on the CAMCE discussed in the main text (Section 5.4) as this quantity is comparable across models with different $K$. For each individual $i$, we calculate our estimate of CAMCE using their moderators $\bm{X}_{i}$ and compare this against the true value, which can be calculated by plugging in the true values of $\pi_{k}(\bm{X}_{i})$ and $\delta_{jk}(l,l^{\prime})$ into Equation (11). We run models with $K\in\{2,3,4\}$ with both split-sample and full data methods discussed above.
@@
Figure A10 reports the ratio of the between-group variability over the total variability across the 1,000 simulations for $K\in\{2,3,4\}$. With $K=2$, we already able to explain 50% of the variability in the data. As expected, $K=2$ shows considerably lower $B_{K}/T$ than higher $K$’s, suggesting its groups are less distinct—or, equivalently, more internally heterogeneous—than $K\in\{3,4\}$. There is limited improvement in quality with $K=4$, which is consistent with the earlier results that the correct choice ($K=3$) adequately summarizes the variability in the data.
@@
Supplementary Material K Additional Results for Immigration Conjoint Experiment
@@
Second, as noted in the main text, we found that sample splitting and refitting the model (see Appendix J.2) was somewhat unstable given different splits of the data. To illustrate this point, Figure A14 shows the 25th-75th percentile (and median) of the AMCEs estimated across twenty repetitions of splitting the data into halves and then using the refitting procedure described above. We address the problem of label switching using a permutation of labels that minimizes the average mean absolute error between all pairs of estimates; we find a permutation by randomly permuting the labels for a randomly chosen set of estimates and repeat this repeatedly until the average mean absolute error stabilizes.
While Figure A14 shows instability in some of the estimated AMCE, it broadly shows a similar result to that in the main text. For example, one group (Group 2 when $K=2$; Group 3 when $K=3$) shows a clear effect of country across most splits whereas one group (Group 1 when $K=2$ and Group 2 when $K=2$) generally shows a large penalty for immigrants who entered without legal authorization.
@@
Finally, Figure 5 in the main text reports the average effect of changing some moderator from $x_{0}$ to $x_{1}$ on $\pi_{k}$, i.e.,

### v5

@@ G.2.2 Updating Moderator Parameters
where $\pi_{k}(\bm{X}_{i})$ and $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ are functions of $\bm{\phi}_{k}$. Note that if $\gamma=0$, this simplifies to a multinomial logistic regression with $\{\mathbb{E}[\bm{1}\{Z_{i}=k\}]\}_{k=1}^{K}$ as the outcome. We perform the $M$-Step using a standard optimizer (e.g., L-BFGS) to optimize $Q_{\phi}$ and thus obtain $\bm{\phi}^{(t+1)}$.
G.3 Classification Maximum Likelihood
If classification maximum likelihood approach is desired, despite statistical concerns about this procedure’s asymptotic bias (e.g., Bryant and Williamson 1978), it can be easily implemented by adapting the preceeding EM algorithm. Celeux and Govaert (1992) propose the “classification EM” algorithm in the spirit of how $k$-means classification is commonly implemented.
The adjustment proceeds as follows (Celeux and Govaert, 1992, p. 319): after conducting an $E$-step and obtaining $\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\bm{\theta})=\tilde{\pi}_{ik}=p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ for use in evaluating $Q_{\beta}$ and $Q_{\phi}$, perform a classification or “hard assignment”. That is, find $k^{*}_{i}=\operatornamewithlimits{argmax}_{k}\tilde{\pi}_{ik}$, i.e., the most probable cluster for observation $i$ given its observed data and $\bm{\theta}$. In the subsequent $M$-step, use a modified weight $c_{ik}=1$ if $k=k^{*}_{i}$ and otherwise $c_{ik}=0$ in lieu of $\tilde{\pi}_{ik}$.
G.4 Degrees of Freedom
@@
G.5 Computational Improvements
While the algorithm above provides a valid way to locate a posterior mode, our estimation problem is complex and high-dimensional. Furthermore, given the complex posterior implied by mixture of experts models, we derived a number of computational strategies to improve convergence. We use the SQUAREM algorithm (Varadhan and Roland 2008). Our software provides the option to use a generalized EM algorithm to update $\bm{\beta}$ using a conjugate gradient approach and $\bm{\phi}$ using a few steps of L-BFGS.
@@ G.5 Computational Improvements
Supplementary Material H Extensions to the Basic Model
@@ H.1 Repeated Observations
H.2 Forced Choice Conjoint Design
@@ H.3 Standardization Weights
This has no implication on the rank of the stacked $\bm{F}_{g}$ (and thus the results in Appendix F) as they are all positive and thus only slightly modify the $E$-Step.
We employ weights in all of our analyses to account for the fact that different factors $j$ may have different number of levels $L_{j}$. We use a generalization of the weights in Bondell and Reich (2009) to the case of penalized differences. Specifically, consider the over-parameterized model in Appendix H.4 where the penalty can be written entirely on the differences $\bm{\delta}_{\mathrm{Main}}$, $\bm{\delta}_{\mathrm{Int}}$, $\bm{\delta}_{\mathrm{Main-Copy}}$. Note that each of those penalties has a simple (group) LASSO form and thus we adopt the approach in Lim and Hastie (2015) of weighting by the Frobenius norm of the associated columns in $\bm{T}_{\mathrm{LOG}}$, i.e. the over-parameterized design matrix. At slight abuse of notation, define $[\bm{T}_{\mathrm{LOG}}]_{g}$ as the columns of $\bm{T}_{\mathrm{LOG}}$ corresponding to the differences penalized in the (group) lasso $g$, the weight can be expressed as follows:
@@
H.4 Latent Overlapping Groups
@@
Supplementary Material I Estimators
@@
I.1 Factorial designs
I.1.1 Without restrictions on randomization
@@
I.1.2 With restrictions on randomization
@@
I.2 Conjoint designs
I.2.1 Without restrictions on randomization
@@
I.2.2 With restrictions on randomization
Similar to Appendix I.1.2, adjustments to estimation need to be made when we have restricted randomizations. We again will do this by dropping profiles that have levels of factors not allowable for all levels of the factor(s) whose effects we are estimating (e.g., profiles with “4th grade” for education when estimating an effect for profession). However, now we estimate the effect for the right profile and the effect for the left profile, and then average the two (they should be equal under symmetry). When estimating the effect for the right profile, therefore, we will only drop pairings if the right profile has a level that is not allowed for some level of the factor we are estimating an effect of. For example, dropping pairings where the right profile has “4th grade” as level of education when estimating main effects of profession because “doctor” cannot have level “4th grade.” Again, this will drop more profiles than those that are not allowed under randomization to ensure an “apples-to-apples” comparison across levels of profession.
@@ I.2.2 With restrictions on randomization
Supplementary Material J Quantification of Uncertainty
@@
$\mathcal{I}_{L}(\hat{\bm{\theta}})=E_{p\left(\{Z_{i}\}_{i=1}^{N}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N},\hat{\bm{\theta}}\right)}\left[-\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}\bm{\theta}^{\top}}\right]-\mathrm{Var}_{p\left(\{Z_{i}\}_{i=1}^{N}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N},\hat{\bm{\theta}}\right)}\left[\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}}\right]$ (A2)
@@
$\begin{split}\tilde{S}^{c}(\bm{\phi}_{k})&=\sum_{i=1}^{N}\left[z_{ik}-\pi_{ik}\right]\bm{X}_{i}+\frac{\partial\log p(\{\bm{\phi}_{k}\})}{\partial\bm{\phi}_{k}}+\\ &\sum_{k^{\prime}=1}^{K}m\gamma\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}}-\lambda\gamma\bar{\pi}_{k^{\prime}}^{\gamma-1}\cdot\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\cdot\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]\end{split}$
@@
$\sum_{k^{\prime}=1}^{K}-\lambda\gamma$ $\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]\left[I(\gamma\notin\{0,1\})\cdot(\gamma-1)\bar{\pi}_{k^{\prime}}^{\gamma-2}\cdot{\color[rgb]{0,0,0}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\right]}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{\ell}}\right]^{\top}+\bar{\pi}_{k^{\prime}}^{\gamma-1}\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}\right]$
@@
$\mathrm{Cov}\left[\tilde{S}^{c}(\mu),\tilde{S}^{c}(\mu)\right]$ $=$
$\sum_{i=1}^{N}\Bigg{[}\sum_{k=1}^{K}\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})$ $\left(I[k=k^{\prime}]-\mathbb{E}(z_{ik^{\prime}})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}^{\top}\right]\Bigg{]}$
@@
J.3 Standard Errors on Other Quantities of Interest
Given the above results, we derive an approximate covariance matrix on $\hat{\bm{\theta}}$. We calculate uncertainty on other quantities of interest, e.g. AMCE and marginal effects, using the multivariate delta method. As almost all of our quantities of interest can be expressed as (weighted) sums or averages over individuals $i\in\{1,\cdots,N\}$, calculating the requisite gradient for the multivariate delta method simply requires calculating the relevant derivative for each observation. For example, all derivatives needed in the AMCE are of the following form; see Appendix I for more details.
@@
Supplementary Material K Simulations
@@
K.1 Setup
We generate the $\bm{\beta}_{k}$ used in our simulations following Equation 3.5 and calibrating their implied AMCEs to be roughly comparable to the magnitude found in our empirical example, i.e. ranging between around $-0.30$ and $0.30$. The $\bm{\beta}_{k}$ and $\{\bm{\phi}_{k}\}_{k=2}^{3}$ used in all simulations are determined using one draw from the following procedure:
@@
In terms of simulating the treatment profiles and outcome, for each individual $i$, we draw a group membership $Z_{i}$ using $\bm{\pi}_{i}$ generating using $\bm{X}_{i}$, $\bm{\phi}$ and Equation 3.5. For each task $t$, we then randomly draw a pair of treatments and then, given $Z_{i}$, draw the outcome $Y_{i}$ given their observed treatments using the model in the main text.
@@
To address this issue, we consider an alternative procedure based on sample splitting. We fit the model using half of the data (selected at random) and then refit the model. To refit the model, we hold fixed the sparsity pattern estimated in the original estimation hold (i.e., which levels are fused together) using a tolerance of $10^{-3}$. We also fix the estimated moderator relationship, i.e. $\pi_{k}(\bm{X}_{i})$, and only estimate the treatment effect coefficients after fusion. Algorithm A2 states the procedure. To calculate the average marginal effects, as noted in Appendix I, we use the empirical distribution of treatments to marginalize over other factors. In this split version, we also use the distribution from the full dataset.
@@
K.3.1 Data-Driven Choice of $K$
First, as noted in the main text, a common approach to choosing $K$ can be information criterion. We use the BIC to calibrate our choice of $\lambda$, i.e. pick the $\lambda$ that minimizes the BIC. In our simulations, we compare the BIC across $K\in\{1,2,3,4\}$ to see which it would suggest choosing. Table K.3.1 reports the probability of each $K$ being chosen across 1,000 simulations. It shows that, even for the smallest data size, the BIC correctly identifies $K=3$. The probability of correct selection rises as the sample size grows. However, as we note in the main text, this simulation example has relatively well separated clusters, and correctly specified likelihoods, and thus the information criterion approach is expected to perform well.
Sample Size $K=1$ $K=2$ $K=3$ $K=4$
1,000 People (5 Tasks) 0 0.01 0.941 0.049
2,000 People (10 Tasks) 0 0.00 0.999 0.001
4,000 People (10 Tasks) 0 0.00 0.994 0.006
Table A2: Probability of $K$ being chosen using smallest BIC
Other criterion based on cross-validation—e.g., splitting the sample and taking the model with the highest out-of-sample predictive likelihood or lowest RMSE—also show a high probability of choosing $K=3$ (84% for the smallest sample size and 97-98% for the larger sample sizes).
K.3.2 Effect of Choice of $K$ on Estimates
We first consider how different choices of $K$ impact our results in the simulation study. To do this, we focus on the CAMCE discussed in the main text (Section 5.4) as this quantity is comparable across models with different $K$. For each individual $i$, we calculate our estimate of CAMCE using their moderators $\bm{X}_{i}$ and compare this against the true value, which can be calculated by plugging in the true values of $\pi_{k}(\bm{X}_{i})$ and $\delta_{jk}(l,l^{\prime})$ into Equation (5.2). We run models with $K\in\{2,3,4\}$ with both split-sample and full data methods discussed above.
@@
Figure A10 reports the ratio of the between-group variability over the total variability across the 1,000 simulations for $K\in\{2,3,4\}$. With $K=2$, we already able to explain around 50% of the variability in the data. As expected, $K=2$ shows considerably lower $B_{K}/T$ than higher $K$’s, suggesting its groups are less distinct—or, equivalently, more internally heterogeneous—than $K\in\{3,4\}$. There is limited improvement in quality with $K=4$, which is consistent with the earlier results that the correct choice ($K=3$) adequately summarizes the variability in the data.
@@
Supplementary Material L Additional Results for Immigration Conjoint Experiment
@@
Second, as noted in the main text, we found that sample splitting and refitting the model (see Appendix K.2) was somewhat unstable given different splits of the data. To illustrate this point, Figure A14 shows the 25th-75th percentile (and median) of the AMCEs estimated across twenty repetitions of splitting the data into halves and then using the refitting procedure described above. We address the problem of label switching using a permutation of labels that minimizes the average mean absolute error between all pairs of estimates; we find a permutation by randomly permuting the labels for a randomly chosen set of estimates and repeat this repeatedly until the average mean absolute error stabilizes.
While Figure A14 shows instability in some of the estimated AMCE, it broadly shows a similar result to that in the main text. For example, one group (Group 2 when $K=2$; Group 3 when $K=3$) shows a clear effect of country across most splits whereas one group (Group 1 when $K=2$ and Groups 1 and 2 when $K=2$) generally shows a large penalty for immigrants who entered without legal authorization.
@@
Third, Figure 6 in the main text reports the average effect of changing some moderator from $x_{0}$ to $x_{1}$ on $\pi_{k}$, i.e.,
@@
Next, we discuss the two-factor interactions. The largest average marginal interaction effect (AMIE) was found between education and job in the three group analysis. This is visualized in Figure A16. The largest AMIE occurs between the levels of Teacher and High School and has magnitude of 0.0021.
Compared in magnitude to the AME, which for education was on average 0.111 and for job was on average 0.0237, this is clearly negligible. Given this, we have little hope of finding substantial higher-order interactions in this example.
If higher-order interactions were of interest, a pre-processing step to do some basic screening (see, e.g., Shi, Wang and Ding, 2023) might be implemented on the full dataset to a priori reduce the number of interactions considered. The sparsity inducing penalties of our method would then impose additional regularization.
Figure A16: The average marginal interaction effect between education and job.
Finally, we briefly remark upon choosing $K$ using an information criterion. While this works well in the simulated example (see Appendix K.3.1), we find less clear results on the full data. Table K.3.1 the results of optimizing the BIC over $\lambda$ for $K\in\{1,2,3,4\}$ as well as optimizing the AIC over $\lambda$. It shows that, if one uses the BIC, this suggests $K=1$. However, if one uses the AIC, this suggests $K=4$.
Table A3: Information criterion for different $K$